OpenStudy (anonymous):

Solve: 3((sec^2)x)-4 =0

OpenStudy (anonymous):

\[3\sec ^{2}x-4=0\] This is a better view of the equation.

OpenStudy (anonymous):

\[\sec^2(x)=\frac{4}{3}\] is a start, so \[\cos^2(x)=\frac{3}{4}\]

OpenStudy (anonymous):

that makes \[\cos(x)=\pm\frac{\sqrt3}{2}\] and now you are mostly done, so long as you know a number whose cosine is \(\frac{\sqrt3}{2}\) and another number whose cosine is \(-\frac{\sqrt3}{2}\)

OpenStudy (anonymous):

How do you know if it is in quadrant I, II, III, or IV?

OpenStudy (mertsj):

cosine is like x. If it is positive, it is in quadrant I, or IV If it is negative, it is in quadrant II or III

OpenStudy (mertsj):

So if your domain is 0 to 360, there should be 4 answers...1 in each quadrant.